OpenStudy (anonymous):

i have an equation in algebra i need help with...it involves square roots with variables and exponents...anyone?

OpenStudy (saifoo.khan):

Sure.

OpenStudy (saifoo.khan):

i can try.

OpenStudy (anonymous):

the equation is \[\sqrt{14x}\times \sqrt{2x}\]

OpenStudy (anonymous):

thanksss(:!

OpenStudy (anonymous):

Sqrt[14x * 2x]

OpenStudy (saifoo.khan):

then: root 28x^2

OpenStudy (anonymous):

THANK YOUUUUUU